# Anisha Banerjee1, Yonatan Yehezkeally2, Antonia Wachter-Zeh1, and Eitan Yaakobi3

### v1

@@ I Introduction
Initial work in this area focused on devising accurate mathematical models for the sequencer or formulating efficient error-correcting codes that incorporate these models. Notably, [maoModels2018, hulettCoding2021, mcbainFiniteState2022a, mcbainInformationRatesNoisy2024] examined the channel from an information-theoretic perspective in an effort to understand its capacity. In particular, the authors of [maoModels2018] proposed a channel model that considers ISI, deletions, and random measurement noise and also derived suitable upper bounds for the capacity of this channel. On the contrary, [hulettCoding2021] used a more deterministic model and developed an algorithm to calculate its capacity. The authors of [mcbainFiniteState2022a, mcbainInformationRatesNoisy2024] studied a finite-state semi-Markov channel (FSMC)–based model for nanopore sequencing that accounts for ISI, duplications, and noisy measurements. They estimated the achievable information rates of this noisy nanopore channel by formulating efficient algorithms to perform this computation. Another exciting direction seeks to facilitate the accurate decoding of DNA fragments, despite sample duplications and background noise, by designing codes directly based on the current signals produced by the nanopore for each sequence of nucleotides [vidalConcatenatedNanoporeDNA2024, vidalErrorBoundsDecoding2023, vidalUnionBoundGeneralized2023].
@@
A subset of prior work [banerjeeCorrectingSingleDeletion2024a, banerjeeErrorCorrectingCodesNanopore2024, cheeCodingSchemeNoisy2024, sunBoundsConstructionsEll2024] studied a specific channel model of the nanopore sequencer inspired by [maoModels2018]. In particular, these works considered the nanopore sequencer to be the concatenation of three distinct channels, as depicted in Fig. 1. The first component that emulates the ISI effect is parameterized by $\ell$ and demonstrates how the observed current is influenced by the $\ell>1$ consecutive nucleotides present in the pore at any moment. This stage is conceptualized as a window of length $\ell$ that slides over an input sequence, shifting by a single position after each time step. This produces a sequence of $\ell$-mers, or a string of $\ell$ symbols, which is then fed to a discrete memoryless channel (DMC) that transforms each $\ell$-mer into a discrete voltage level based on a deterministic function, assumed to be the Hamming weight when the input is binary, and the composition function otherwise. The final stage introduces errors, such as substitutions, deletions, or duplications, that corrupt the sequence of discrete voltage levels. While [banerjeeCorrectingSingleDeletion2024a, banerjeeErrorCorrectingCodesNanopore2024] assumed that the error component introduces either at most one substitution or at most one deletion in the final channel output, [sunBoundsConstructionsEll2024] considered multiple substitutions and investigated bounds and constructions of suitable substitution-correcting codes for this channel, when $\ell=2$. In contrast, [cheeCodingSchemeNoisy2024] proposed constrained codes for this channel to combat duplication and deletion errors. This model is also similar to the transverse-read channel [cheeTransverseReadCodesDomainWall2023, yerushalmiCapacityWeightedRead2024], which is important for racetrack memories. We now state the problem formally.
[…]
@@
The primary contribution of this work is to demonstrate that for $\ell\geqslant 2$, the minimum redundancy required for any code of length $n$ that can correct $t\geqslant 2$ substitutions in $\ell$-read vectors is at least $t\log_{2}n-O(1)$ bits (Theorem 4), which starkly contrasts with the minimum redundancy bound of $\log_{2}\log_{2}n-o(1)$ [banerjeeErrorCorrectingCodesNanopore2024] for the case of $t=1$. Following the introduction of essential notations and definitions in Section II, we present the proof of this lower bound, based on the clique cover technique [knuthSandwich1994, chrisnataCorrecting2022], in Section III. With this unfortunate result, we show in Construction A that a naive construction achieves this bound up to a constant.
@@ II Preliminaries
For any $q\geqslant 2$, we let $\Sigma_{q}$ represent the $q$-ary alphabet $\{0,1,\ldots,q-1\}$. The set of all $q$-ary sequences of length $n$ is indicated by $\Sigma_{q}^{n}$, with $\Sigma_{q}^{0}$ meant to indicate the empty set. Additionally, we let $\Sigma_{q}^{\geqslant n}=\cup_{i=n}^{\infty}\Sigma_{q}^{i}$. For any two integers $a,b$ such that $a\leqslant b$, $[a,b]$ is used to denote the set $\{a,a+1,\ldots,b\}$, and $[b]\triangleq[1,b]$ for $b\geqslant 1$. We also use the following notation to denote element-wise modulo operation on a vector ${\boldsymbol{y}}\in\Sigma_{q}^{n}$.
@@
Throughout this work, we focus on ${q=2}$, and the output of the channel we consider is a sequence of readings of a sliding window moved across ${\boldsymbol{x}}$, as defined below.
@@
[banerjeeCorrectingSingleDeletion2024a, banerjeeErrorCorrectingCodesNanopore2023] The $\ell$-read vector of any ${{\boldsymbol{x}}\in{{\Sigma}}_{\!2}^{n}}$ is of length $n+\ell-1$ over $\Sigma_{\ell+1}$, and is defined as
@@
The $i$-th element of $\mathcal{R}_{\ell}({\boldsymbol{x}})$ is denoted as $\mathcal{R}_{\ell}({\boldsymbol{x}})_{i}$; that is, $\mathcal{R}_{\ell}(\boldsymbol{x})_{i}=\operatorname{wt}({\boldsymbol{x}}_{i-\ell+1}^{i})$. We will omit $\ell$ from the subscript of $\mathcal{R}_{\ell}({\boldsymbol{x}})$ whenever it is clear from the context. It is worth pointing out that Definition 1 can be extended to the non-binary alphabet, by considering compositions instead of Hamming weights [banerjeeErrorCorrectingCodesNanopore2024]. The composition of a vector $q$-ary ${\boldsymbol{x}}$ refers to the count of each symbol in $\Sigma_{q}$ as it appears in ${\boldsymbol{x}}$.
[…]
@@
As mentioned previously, [cheeTransverseReadCodesDomainWall2023, yerushalmiCapacityWeightedRead2024] examine a similar model, wherein the output sequence, called the transverse-read vector, is essentially a substring of the $\ell$-read vector for specific parameter choices.
Since we are interested in codes that correct up to $t$ substitutions in $\ell$-read vectors, it is essential to precisely define what is meant by an error-correcting code in the framework of our channel. Similarly to [banerjeeErrorCorrectingCodesNanopore2024], a code is said to be a $t$-substitution $\ell$-read code if for any distinct ${\boldsymbol{x}},{\boldsymbol{y}}$ from this code, it holds that $d_{H}(\mathcal{R}({\boldsymbol{x}}),\mathcal{R}({\boldsymbol{y}}))>2t$.
@@ III Correcting Multiple Substitutions
This section aims to establish an upper bound on the size of a code that corrects $t$ substitutions in $\ell$-read vectors, where $t\geqslant 2$ is constant; the case $t=1$ was thoroughly analysed in [banerjeeErrorCorrectingCodesNanopore2024]. To accomplish this, we apply the clique cover technique, which was also used in [chrisnataCorrecting2022, banerjeeErrorCorrectingCodesNanopore2024] for the case of $t=1$. This method considers a graph $\mathcal{G}(n)$ that contains vertices corresponding to all vectors in $\Sigma^{n}_{2}$. In this graph, any two vertices representing distinct binary vectors, say ${\boldsymbol{x}}$ and ${\boldsymbol{y}}$ in $\Sigma^{n}_{2}$, are considered adjacent if and only they satisfy $d_{H}(\mathcal{R}({\boldsymbol{x}}),\mathcal{R}({\boldsymbol{y}}))\leqslant 2t$. Consequently, any subset of vertices in $\mathcal{G}(n)$ where no two vertices are adjacent (namely, an independent set) constitutes a $t$-substitution $\ell$-read code. In contrast, a clique in the graph is a set of vertices that are all pair-wise adjacent. The formal definition of a clique cover is stated below.
[…]
@@
From [chrisnataOptimal2020a, knuthSandwich1994], the following result is widely known.
@@
[…]
[banerjeeErrorCorrectingCodesNanopore2023, Definition 3] For a positive integer $p$, define a permutation $\pi_{p}$ on ${{\Sigma}}_{\!2}^{n}$ as follows. For all ${\boldsymbol{x}}\in{{\Sigma}}_{\!2}^{n}$, arrange the coordinates of ${\boldsymbol{x}}_{1}^{p\ell\lfloor n/(p\ell)\rfloor}$ in a matrix $X\in\Sigma^{p\lfloor n/(p\ell)\rfloor\times\ell}$, by row (first fill the first row from left to right, then the next, etc.). Next, partition $X$ into sub-matrices of dimension $p\times 2$ (if $\ell$ is odd, we ignore $X$’s right-most column). Finally, going through each sub-matrix (from left to right, and then top to bottom), we concatenate its rows, to obtain $\pi_{p}({\boldsymbol{x}})$ (where unused coordinates from ${\boldsymbol{x}}$ are appended arbitrarily).
@@
The subsequent definition presents the core component of our clique cover, and borrows ideas from [chrisnataCorrecting2022, banerjeeErrorCorrectingCodesNanopore2024].
[…]
@@
$\Lambda^{(1)}_{p}$ =$=\left\{(01)^{j}(10)^{p-j}:j\in[p]\right\},$
$\Lambda^{(2)}_{p}$ =$=\left\{(10)^{j}(01)^{p-j}:j\in[p]\right\}$
@@
[…]
@@
[…]
@@
$\mathcal{Q}(m,p)$ =$=\left\{\left\{{\boldsymbol{x}}\right\}:{\boldsymbol{x}}\in\widetilde{\Lambda}_{p}^{m}\right\}\cup\left\{Q_{\gamma}^{(0)},Q_{\gamma}^{(1)}:\gamma\in\Gamma_{1}\right\}\cup\cdots$
@@
To specify the set of indices in $[2pm+\ell-1]$ at which $\mathcal{R}({\boldsymbol{x}})$ and $\mathcal{R}({\boldsymbol{y}})$ disagree, we let the index at which the substring $({\boldsymbol{\alpha}}_{h})^{m_{h}}$ (or $({\boldsymbol{\beta}}_{h})^{m_{h}}$) starts in $\pi_{p}({\boldsymbol{x}})$ (or $\pi_{p}({\boldsymbol{y}})$) be given by $c_{h}$, for all $h\in[s]$. Thus, the starting and ending indices of each $({\boldsymbol{\alpha}}_{h})^{m_{h}}$ (or $({\boldsymbol{\beta}}_{h})^{m_{h}}$) in $\pi_{p}({\boldsymbol{x}})$ (or $\pi_{p}({\boldsymbol{y}})$) are $c_{h}$ and $c_{h}+2m_{h}-1$, which map to the positions $f_{\pi}(c_{h})$ and $f_{\pi}(c_{h}+2m_{h}-1)$ in ${\boldsymbol{x}}$ (or ${\boldsymbol{y}}$), respectively. Since for all $i\in\bigcup_{h\in[s]}\{f_{\pi}(c_{h}),\ldots,f_{\pi}(c_{h}+2m_{h}-1)\}$, $x_{i}\neq y_{i}$, it is possible that for certain instances of $r$, $\mathcal{R}(\boldsymbol{{\boldsymbol{x}}})_{r+\ell-1}\neq\mathcal{R}(\boldsymbol{{\boldsymbol{y}}})_{r+\ell-1}$.
@@
is also a clique (the case $i=1$ was already proven in [banerjeeErrorCorrectingCodesNanopore2023]). It now remains to show that $\pi_{p}({\boldsymbol{x}})$ belongs to at least one clique in $\mathcal{Q}(m,p)$ for any ${\boldsymbol{x}}\in{{\Sigma}}_{\!2}^{2pm}$. For simplicity, we use the fact that $\pi_{p}$ is a permutation to show instead that any such ${\boldsymbol{x}}$ is itself a member.
To this end, consider a ${\boldsymbol{x}}\in{{\Sigma}}_{\!2}^{2pm}$ and the set =$\mathcal{I}=\{i_{1},\ldots,i_{|\mathcal{I}|}\}$ where $i_{1}<\cdots<i_{|\mathcal{I}|}$, that satisfies $i_{h}-1\equiv 0\pmod{2p}$ and $\pi_{p}({\boldsymbol{x}})_{i_{h}}^{i_{h}+2p-1}\in\Lambda_{p}$ for all $h\in[|\mathcal{I}|]$. We consider $\mathcal{I}$ to be exhaustive, i.e., there exists no $i\in[2p(m-1)+1]$ such that $i-1\equiv 0\pmod{2p}$, $\pi_{p}({\boldsymbol{x}})_{i}^{i+2p-1}\in\Lambda_{p}$ and $i\not\in\mathcal{I}$.
@@
It follows from Section III that
@@
[…]
For $s\geqslant 0$, $p=\lceil\frac{1}{2}(1-\epsilon)\log_{2}n\rceil$ and $10<\lambda<$ 1,
@@
[…]
For $10<\lambda<$ 1 and a positive integer $t$, it holds that
@@
We apply the following inequality [bar-levAdversarialTornPaperCodes2023]
$\log_{2}\binom{u+v}{u}$$\leqslant$ $u(2\log(e)+\log(\frac{v}{u})),$
@@
$\leqslant$ $\frac{t-1}{\log(1/\lambda)}\left\lparen\frac{t-1}{\log(1/\lambda)}\lambda^{-1/\log\lambda}\right\rparen^{t-1}$
$+\lim_{n\to\infty}\sum_{r=1+\frac{t-1}{\log(1/\lambda)}}^{n}\left\lparen\frac{2(t-1)}{\log(1/\lambda)}\lambda^{-1/\log\lambda}\right\rparen^{t-1}\lambda^{r/2}$
$=$ $\frac{t-1}{\log(1/\lambda)}+\left\lparen\frac{2(t-1)}{e\log(1/\lambda)}\right\rparen^{t-1}\frac{\lambda^{{1/2-(t-1)/(2\log\lambda})}}{1-\lambda^{1/2}}$
$=$ ($\left\lparen\frac{t-1}{e\log(1/\lambda)}\right\rparen^{t-1}\left\lparen\frac{t-1}{\log(1/\lambda)}+2^{t-1}\frac{\lambda^{1/2}e^{-(t-1)/2}}{1-\lambda^{1/2}}\right\rparen.$ ) .
@@
This theorem suggests that for $\ell\geqslant 2$ and $t\geqslant 2$, the minimum redundancy required by any $t$-substitution-correcting code for a channel that produces $\ell$-read vectors is, up to a fixed addend, the same as that of the classical substitution channel [rothIntroductionCodingTheory2007, Theorem 4.3, Lemma 4.8], assuming $t$ is fixed with respect to $n$. For $\ell=2$, this result is also proved in [sunBoundsConstructionsEll2024]. Theorem 4 is dispiriting in light of [banerjeeErrorCorrectingCodesNanopore2024, Lemma 14], which shows that to correct a single ($t=1$) substitution in $\ell$-read vectors, when $\ell\geqslant 3$, $\log_{2}\log_{2}n+o(1)$ redundant bits are sufficient.
Next, we focus on designing a suitable error-correcting code by leveraging the fact that ${\boldsymbol{x}}$ can be directly inferred from the first or last $n$ elements of $\mathcal{R}({\boldsymbol{x}})\bmod 2$ [banerjeeErrorCorrectingCodesNanopore2023, Proposition 1]. This leads us to the following naive $t$-substitution $\ell$-read code construction is optimal up to a constant.
@@
Evidently, Construction A requires $t\log_{2}n$ redundant bits. Thus, for $t\geqslant 2$ and $\ell\geqslant 2$, it is a $t$-substitution $\ell$-read code that is optimal up to a constant.

### v2

@@ I Introduction
Initial work in this area focused on devising accurate mathematical models for the sequencer or formulating efficient error-correcting codes that incorporate these models. Notably, [1, 2, 3, 4] examined the channel from an information-theoretic perspective in an effort to understand its capacity. In particular, the authors of [1] proposed a channel model that considers ISI, deletions, and random measurement noise and also derived suitable upper bounds for the capacity of this channel. On the contrary, [2] used a more deterministic model and developed an algorithm to calculate its capacity. The authors of [3, 4] studied a finite-state semi-Markov channel (FSMC)–based model for nanopore sequencing that accounts for ISI, duplications, and noisy measurements. They estimated the achievable information rates of this noisy nanopore channel by formulating efficient algorithms to perform this computation. Another exciting direction seeks to facilitate the accurate decoding of DNA fragments, despite sample duplications and background noise, by designing codes directly based on the current signals produced by the nanopore for each sequence of nucleotides [5, 6, 7].
@@
A subset of prior work [8, 9, 10, 11] studied a specific channel model of the nanopore sequencer inspired by [1]. In particular, these works considered the nanopore sequencer to be the concatenation of three distinct channels, as depicted in Fig. 1. The first component that emulates the ISI effect is parameterized by $\ell$ and demonstrates how the observed current is influenced by the $\ell>1$ consecutive nucleotides present in the pore at any moment. This stage is conceptualized as a window of length $\ell$ that slides over an input sequence, shifting by a single position after each time step. This produces a sequence of $\ell$-mers, or a string of $\ell$ symbols, which is then fed to a discrete memoryless channel (DMC) that transforms each $\ell$-mer into a discrete voltage level based on a deterministic function, assumed to be the Hamming weight when the input is binary, and the composition function otherwise. The final stage introduces errors, such as substitutions, deletions, or duplications, that corrupt the sequence of discrete voltage levels. While [8, 9] assumed that the error component introduces either at most one substitution or at most one deletion in the final channel output, [11] considered multiple substitutions and investigated bounds and constructions of suitable substitution-correcting codes for this channel, when $\ell=2$. In contrast, [10] proposed constrained codes for this channel to combat duplication and deletion errors. This model is also similar to the transverse-read channel [12, 13], which is important for racetrack memories. We now state the problem formally.
Question 1
@@
The primary contribution of this work is to demonstrate that for $\ell\geqslant 2$, the minimum redundancy required for any code of length $n$ that can correct $t\geqslant 2$ substitutions in $\ell$-read vectors is at least $t\log_{2}n-O(1)$ bits (Theorem 17), which starkly contrasts with the minimum redundancy bound of $\log_{2}\log_{2}n-o(1)$ [9] for the case of $t=1$. Following the introduction of essential notations and definitions in Section II, we present the proof of this lower bound, based on the clique cover technique [14, 15], in Section III. With this unfortunate result, we show in Construction˜A that a naive construction achieves this bound up to a constant.
@@ II Preliminaries
For any $q\geqslant 2$, we let $\Sigma_{q}$ represent the $q$-ary alphabet $\{0,1,\ldots,q-1\}$. The set of all $q$-ary sequences of length $n$ is indicated by $\Sigma_{q}^{n}$, with $\Sigma_{q}^{0}$ meant to indicate the empty set. Additionally, we let $\Sigma_{q}^{\geqslant n}=\cup_{i=n}^{\infty}\Sigma_{q}^{i}$. For any two integers $a,b$ such that $a\leqslant b$, $[a,b]$ is used to denote the set $\{a,a+1,\ldots,b\}$, and $b]\triangleq[1,b]$ for $b\geqslant 1$. We also use the following notation to denote element-wise modulo operation on a vector ${\boldsymbol{y}}\in\Sigma_{q}^{n}$.
@@
Throughout this work, we focus on q=2{q=2}, and the output of the channel we consider is a sequence of readings of a sliding window moved across $\boldsymbol{x}$}, as defined below.
@@
[8, 16] The $\ell$-read vector of any ${\boldsymbol{x}}\in{{\Sigma}}_{\!2}^{n}$} is of length $n+\ell-1$ over $\Sigma_{\ell+1}$, and is defined as
@@
The $i$-th element of $\mathcal{R}_{\ell}({\boldsymbol{x}})$ is denoted as $\mathcal{R}_{\ell}({\boldsymbol{x}})_{i}$; that is, $\mathcal{R}_{\ell}(\boldsymbol{x})_{i}=\operatorname{wt}({\boldsymbol{x}}_{i-\ell+1}^{i})$. We will omit $\ell$ from the subscript of $\mathcal{R}_{\ell}({\boldsymbol{x}})$ whenever it is clear from the context. It is worth pointing out that Definition 2 can be extended to the non-binary alphabet, by considering compositions instead of Hamming weights [9]. The composition of a vector $q$-ary $\boldsymbol{x}$} refers to the count of each symbol in $\Sigma_{q}$ as it appears in $\boldsymbol{x}$}.
Example 3
@@
As mentioned previously, [12, 13] examine a similar model, wherein the output sequence, called the transverse-read vector, is essentially a substring of the $\ell$-read vector for specific parameter choices.
Since we are interested in codes that correct up to $t$ substitutions in $\ell$-read vectors, it is essential to precisely define what is meant by an error-correcting code in the framework of our channel. Similarly to [9], a code is said to be a $t$-substitution $\ell$-read code if for any distinct ${\boldsymbol{x}},{\boldsymbol{y}}$ from this code, it holds that $d_{H}(\mathcal{R}({\boldsymbol{x}}),\mathcal{R}({\boldsymbol{y}}))>2t$.
@@ III Correcting Multiple Substitutions
This section aims to establish an upper bound on the size of a code that corrects $t$ substitutions in $\ell$-read vectors, where $t\geqslant 2$ is constant; the case $t=1$ was thoroughly analysed in [9]. To accomplish this, we apply the clique cover technique, which was also used in [15, 9] for the case of $t=1$. This method considers a graph $\mathcal{G}(n)$ that contains vertices corresponding to all vectors in $\Sigma^{n}_{2}$. In this graph, any two vertices representing distinct binary vectors, say $\boldsymbol{x}$} and $\boldsymbol{y}$} in $\Sigma^{n}_{2}$, are considered adjacent if and only they satisfy $d_{H}(\mathcal{R}({\boldsymbol{x}}),\mathcal{R}({\boldsymbol{y}}))\leqslant 2t$. Consequently, any subset of vertices in $\mathcal{G}(n)$ where no two vertices are adjacent (namely, an independent set) constitutes a $t$-substitution $\ell$-read code. In contrast, a clique in the graph is a set of vertices that are all pair-wise adjacent. The formal definition of a clique cover is stated below.
Definition 4
@@
From [17, 14], the following result is widely known.
@@
Definition 6
[16, Definition 3] For a positive integer $p$, define a permutation $\pi_{p}$ on ${{\Sigma}}_{\!2}^{n}$ as follows. For all ${\boldsymbol{x}}\in{{\Sigma}}_{\!2}^{n}$, arrange the coordinates of ${\boldsymbol{x}}_{1}^{p\ell\lfloor n/(p\ell)\rfloor}$ in a matrix $X\in\Sigma^{p\lfloor n/(p\ell)\rfloor\times\ell}$, by row (first fill the first row from left to right, then the next, etc.). Next, partition $X$ into sub-matrices of dimension $p\times 2$ (if $\ell$ is odd, we ignore $X$’s right-most column). Finally, going through each sub-matrix (from left to right, and then top to bottom), we concatenate its rows, to obtain $\pi_{p}({\boldsymbol{x}})$ (where unused coordinates from $\boldsymbol{x}$} are appended arbitrarily).
@@
The subsequent definition presents the core component of our clique cover, and borrows ideas from [15, 9].
Definition 9
@@
$\Lambda^{(1)}_{p}$ $=\left\{(01)^{j}(10)^{p-j}:j\in[p]\right\},$
$\Lambda^{(2)}_{p}$ $=\left\{(10)^{j}(01)^{p-j}:j\in[p]\right\}$
@@
$\vdots$ $\qquad\qquad\qquad\qquad\vdots$
@@
$\vdots$ $\qquad\qquad\qquad\qquad\vdots$
@@
$\mathcal{Q}(m,p)$ $=\left\{\left\{{\boldsymbol{x}}\right\}:{\boldsymbol{x}}\in\widetilde{\Lambda}_{p}^{m}\right\}\cup\left\{Q_{\gamma}^{(0)},Q_{\gamma}^{(1)}:\gamma\in\Gamma_{1}\right\}\cup\cdots$
@@
To specify the set of indices in $2pm+\ell-1]$ at which $\mathcal{R}({\boldsymbol{x}})$ and $\mathcal{R}({\boldsymbol{y}})$ disagree, we let the index at which the substring $({\boldsymbol{\alpha}}_{h})^{m_{h}}$ (or $({\boldsymbol{\beta}}_{h})^{m_{h}}$) starts in $\pi_{p}({\boldsymbol{x}})$ (or $\pi_{p}({\boldsymbol{y}})$) be given by $c_{h}$, for all $h\in[s]$. Thus, the starting and ending indices of each $({\boldsymbol{\alpha}}_{h})^{m_{h}}$ (or $({\boldsymbol{\beta}}_{h})^{m_{h}}$) in $\pi_{p}({\boldsymbol{x}})$ (or $\pi_{p}({\boldsymbol{y}})$) are $c_{h}$ and $c_{h}+2m_{h}-1$, which map to the positions $f_{\pi}(c_{h})$ and $f_{\pi}(c_{h}+2m_{h}-1)$ in $\boldsymbol{x}$} (or $\boldsymbol{y}$}), respectively. Since for all $i\in\bigcup_{h\in[s]}\{f_{\pi}(c_{h}),\ldots,f_{\pi}(c_{h}+2m_{h}-1)\}$, $x_{i}\neq y_{i}$, it is possible that for certain instances of $r$, $\mathcal{R}(\boldsymbol{{\boldsymbol{x}}})_{r+\ell-1}\neq\mathcal{R}(\boldsymbol{{\boldsymbol{y}}})_{r+\ell-1}$.
@@
is also a clique (the case $i=1$ was already proven in [16]). It now remains to show that $\pi_{p}({\boldsymbol{x}})$ belongs to at least one clique in $\mathcal{Q}(m,p)$ for any ${\boldsymbol{x}}\in{{\Sigma}}_{\!2}^{2pm}$. For simplicity, we use the fact that $\pi_{p}$ is a permutation to show instead that any such $\boldsymbol{x}$} is itself a member.
To this end, consider a ${\boldsymbol{x}}\in{{\Sigma}}_{\!2}^{2pm}$ and the set $\mathcal{I}=\{i_{1},\ldots,i_{|\mathcal{I}|}\}$ where $i_{1}<\cdots<i_{|\mathcal{I}|}$, that satisfies $i_{h}-1\equiv 0\pmod{2p}$ and $\pi_{p}({\boldsymbol{x}})_{i_{h}}^{i_{h}+2p-1}\in\Lambda_{p}$ for all $h\in[|\mathcal{I}|]$. We consider $\mathcal{I}$ to be exhaustive, i.e., there exists no $i\in[2p(m-1)+1]$ such that $i-1\equiv 0\pmod{2p}$, $\pi_{p}({\boldsymbol{x}})_{i}^{i+2p-1}\in\Lambda_{p}$ and $i\not\in\mathcal{I}$.
@@
It follows from Section˜III that
@@
Lemma 15
For $s\geqslant 0$, $p=\lceil\frac{1}{2}(1-\epsilon)\log_{2}n\rceil$ and $0<\lambda<1$,
@@
Lemma 16
For $0<\lambda<1$ and a positive integer $t$, it holds that
@@
We apply the following inequality [18]
$\log_{2}\binom{u+v}{u} \displaystyle\leqslant$ $u(2\log(e)+\log(\frac{v}{u})),$
@@
[…]
@@
This theorem suggests that for $\ell\geqslant 2$ and $t\geqslant 2$, the minimum redundancy required by any $t$-substitution-correcting code for a channel that produces $\ell$-read vectors is, up to a fixed addend, the same as that of the classical substitution channel [19, Theorem 4.3, Lemma 4.8], assuming $t$ is fixed with respect to $n$. For $\ell=2$, this result is also proved in [11]. Theorem 17 is dispiriting in light of [9, Lemma 14], which shows that to correct a single ($t=1$) substitution in $\ell$-read vectors, when $\ell\geqslant 3$, $\log_{2}\log_{2}n+o(1)$ redundant bits are sufficient.
Next, we focus on designing a suitable error-correcting code by leveraging the fact that $\boldsymbol{x}$} can be directly inferred from the first or last $n$ elements of $\mathcal{R}({\boldsymbol{x}})\bmod 2$ [16, Proposition 1]. This leads us to the following naive $t$-substitution $\ell$-read code construction is optimal up to a constant.
@@
Evidently, Construction˜A requires $t\log_{2}n$ redundant bits. Thus, for $t\geqslant 2$ and $\ell\geqslant 2$, it is a $t$-substitution $\ell$-read code that is optimal up to a constant.